OpenStudy (anonymous):

Find the zeros of the function f(x) = x2 - 7x + 6 by factoring. What is the largest zero of the function? Enter a number only.

OpenStudy (anonymous):

did you try to factor it?

OpenStudy (anonymous):

I'm not sure how I keep getting confused..

OpenStudy (anonymous):

what are the factors of 6?

OpenStudy (anonymous):

3 and 2?

OpenStudy (anonymous):

(x +3)(x +2) =?

OpenStudy (anonymous):

are those the only factors?

OpenStudy (anonymous):

yeah?

OpenStudy (anonymous):

how about -2 and -3?

OpenStudy (anonymous):

(x -3)(x -2) =?

OpenStudy (anonymous):

any others?

OpenStudy (anonymous):

I didn't think about those.. and I don't think so.

OpenStudy (anonymous):

how about 6 and 1? (x +1)(x +6) =?

OpenStudy (anonymous):

might as well finish up -1 and -6 are also factors of 6: (x -1)(x -6) =? which one of those possibilities gives you the equation you are trying to factor?

OpenStudy (anonymous):

would you plug those into the 2 and 7?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

like the -6 and -1 or 3 and 2; would you plug them into the equation to figure it out?

OpenStudy (anonymous):

no, expand those factored forms I gave you; see which one matches the quadratic you are trying to factor.

OpenStudy (anonymous):

(hint: start with the last one I gave you.)

OpenStudy (anonymous):

so -6 factored would be -2 and 3 or -3 and 2 right?

OpenStudy (anonymous):

expand this : (x -1)(x -6)

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

(x -1)(x -6) = x(x -6) + -1(x -6) = ? finish that up, add any like terms, see if it matches your original quadratic